OpenStudy (anonymous):

solve using the quadratic equation. 2x^2-3x=-3

jhonyy9 (jhonyy9):

2x2 -3x+3=0 x_1,_2=(3+/- sqrt(9-24))/4

OpenStudy (anonymous):

2x^2 -3x+3=0 a=2, b=-3, c=3 plug into the quadratic equation which the answer would be: 3+/- i√15 ___________ 4 +/- = plus or minus

OpenStudy (anonymous):

lets move the -3 over 2x^2 - 3x + 3 = 0 you can already see that its impossible to factor since no two numbers can add up to -3 and multiply 3. Thats why you are using the quadratic in this. The general form of this equation is ax^2 + bx + c = 0 and the quadratic formula is (-b +/- sqrt(b^2-4ac))/2a so lets plug in (3 + sqrt(9 - 4(2)(3))/4) and (3 - sqrt(9 - 4(2)(3))/4) then you solve it from there to get the 2 different values of x.

OpenStudy (anonymous):

ok, i got 3+/-sqrt(-15)/4. it says to use i for complex numbers? not sure what that means and if i need to use it here?

jhonyy9 (jhonyy9):

us (3+/- isqrt(15)/4

OpenStudy (anonymous):

thank you!